OpenStudy (anonymous):

HELPP!! Two airplanes leave an airport at the same time. the first airplane flies east at 635mph. the second airplane flies west at 615mph. how long will it take for the airplanes to be 1,000 miles apart?

OpenStudy (anonymous):

Answer = 0.8 hours = 48 minutes Relative Speed of Fast plane = 635 mph + 615 mph = 1250 mph Time to 1000 miles = 1000 miles / 1250 miles per hour = 0.8 hours Speed in opposite direction = v1 + v2 Speed in Same direction = v1-v2 v1 = speed of fast object v2 = speed of slower object

OpenStudy (anonymous):

so 48 minutes?

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

8 hrs?

OpenStudy (anonymous):

its bout flieing so how can it be 48 minutes? :/

OpenStudy (anonymous):

its impossible

OpenStudy (anonymous):

start with the basics : speed = distance/time

OpenStudy (anonymous):

i did

OpenStudy (jhannybean):

Draw yourself a diagram.

OpenStudy (anonymous):

would agree with @Jhannybean

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thats what i got :b

OpenStudy (jhannybean):

@Rachel_K Would you please stop giving answers? It's against the code of conduct here on OS.

OpenStudy (anonymous):

i didn't i explained in the beginning im just confirming what i got

OpenStudy (anonymous):

.

OpenStudy (anonymous):